OpenStudy (anonymous):

Write balanced equation for the following reactions: 1) magnesium carbonate+hydrochloric acid-->magnesium chloride+water+carbon-dioxide, 2) iron+sulphuric acid-->iron(II)sulphate+hydrogen, 3)carbon+oxygen(gas)-->carbon monoxide, 4)iron(III)oxide+carbon-->iron+carbon dioxide, 5) ammonium hydroxide+sulphuric acid-->ammonium sulphate+water

OpenStudy (anonymous):

you know the formulas of all ???

OpenStudy (anonymous):

no

OpenStudy (anonymous):

like for 1 . magnesium carbonate is MgCO3

OpenStudy (anonymous):

no i dont know

OpenStudy (anonymous):

then its very difficult for you . I can't give you direct answers

OpenStudy (anonymous):

you know elements and there symbols ??

OpenStudy (anonymous):

ok can u tell me that how can we calculate the molecular mass of \[Al(OH)_{3}\]

OpenStudy (anonymous):

But I need some other values too

OpenStudy (anonymous):

ops its Molecular mass .Yeah yeah I can :)

OpenStudy (anonymous):

no i just need the formula or molecular mass of \[Al(OH)_{3}\]

OpenStudy (anonymous):

The molar mass is the sum of the masses of each of the atoms in the molecule: Al: 26.98 O: 15.99x3 H: 1.008x3 Total: 77.97 g/mol ~ 78 g/mol (the O and H are multiplied by three because there are three of each in the molecule)

OpenStudy (anonymous):

but i think it is \[(1\times 27)+(3\times 16)+(3\times 1) =78\] is this right

OpenStudy (anonymous):

both are same dear :)

OpenStudy (anonymous):

Is it necessary to write g/mol

OpenStudy (anonymous):

Can u tell me \[Fe _{2}(SO _{4})_{3}\]

OpenStudy (anonymous):

yeah it is necessary

OpenStudy (anonymous):

Fe = 55.845 S = 32.065 O = 15.9994 Therefore: 2*55.845 + 3*32.065 + 12*15.9994 = 399.88 g/mol

OpenStudy (anonymous):

u did Oxygen 12 times because???

OpenStudy (anonymous):

did u multiplied \[4\times 3\]

OpenStudy (anonymous):

(SO4)3 4 x 3 = 12 means 12 no of Oxygen in formula

OpenStudy (anonymous):

did you get it?

OpenStudy (anonymous):

ya thank u

OpenStudy (anonymous):

No problem !!